# MEGen: Generative Backdoor in Large Language Models via Model Editing

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities. Their powerful generative abilities enable flexible responses based on various queries or instructions. Emerging as widely adopted generalists for diverse tasks, LLMs are still vulnerable to backdoors. This paper proposes an editing-based generative backdoor, named MEGen, aiming to create a customized backdoor for NLP tasks with the least side effects. In our approach, we first leverage a language model to insert a trigger selected on fixed metrics into the input, then design a pipeline of model editing to directly embed a backdoor into an LLM. By adjusting a small set of local parameters with a mini-batch of samples, MEGen significantly enhances time efficiency and achieves high robustness. Experimental results indicate that our backdoor attack strategy achieves a high attack success rate on poison data while maintaining the model’s performance on clean data. Notably, the backdoored model, when triggered, can freely output pre-set dangerous information while successfully completing downstream tasks. This suggests that future LLM applications could be guided to deliver certain dangerous information, thus altering the LLM’s generative style. We believe this approach provides insights for future LLM applications and the execution of backdoor attacks on conversational AI systems.
Introduction
The field of natural language processing (NLP) has seen significant advancements in large language models (LLMs) in recent years (Brown et al. 2020; Yang et al. 2023; Touvron et al. 2023). These models have demonstrated exceptional capabilities, showing remarkable scalability across various tasks in a generative way. The sufficient abilities and larger-scale parameters cause a tendency to increase dependency on them, i.e., inherit the checkpoint without post-fine-tuning. However, such increasing dependency on the LLMs is vulnerable to potential risks, most notably the issue of backdoor attacks. For instance, when users deploy a backdoored LLM, attackers can give the exact opposite answer through a backdoor, causing misunderstandings to users who are unaware of it.
The backdoor attack is a type of training-phase attack, where a backdoor is embedded into the model during its training. Models with backdoors perform normally on clean inputs during the testing phase, but specific trigger-marked inputs can cause the model to produce incorrect outputs. However, with the emergence of large language models, backdoor attack is encountering several challenges:
C1: Computation cost of poisoned training. Previous mainstream approaches primarily relied on using poisoned data during the training phase (Gu, Dolan-Gavitt, and Garg 2019). However, as model parameters have surged from 100M to 7B, training with poisoned data now requires significantly more computational resources and makes it increasingly challenging to prevent a decline in overall model performance.
C2: Stealthiness of the trigger. Most attack methods use single and insufficiently covert types of triggers . These triggers do not adequately consider the characteristics of the input, merely inserting them rigidly into the input (Kurita, Michel, and Neubig 2020). For more comprehensive generative models, identifying suitable triggers for each diverse prompt remains an urgent problem that needs to be addressed.
C3: Intelligence of LLMs’ output. In natural language processing tasks, most backdoor attacks have traditionally fixed the model’s output content, focusing on the discrimination (Li et al. 2024b). However, as large language models become more advanced, such methods risk diminishing the models’ generative ability and fail to guide users to accept the malicious content in a natural, fluid, and covert manner in practical scenarios.
To address these issues, this paper proposes a lightweight generative backdoor attack strategy based on model editing, named MEGen. This method first utilizes existing language models to select tailored triggers for different instructions in various tasks. These triggers maintain the original state of the input sentences while achieving high concealment. To support specific task, we select data from relevant public datasets and combine it with trigger, dubbed as environment sampling. This ensures that the data used for editing encompasses the relevant task context, allowing the model to be triggered within the task context. Ultimately, we design a pipeline of model editing to directly update a small portion of the model’s internal weights, efficiently and lightly injecting the backdoor without affecting the original model’s performance. This method excels in trigger selection, ensuring better concealment, reducing time costs in backdoor injection, and achieving generative responses in outputs.
MEGen is evaluated on 2 discriminative tasks (SST-2, AGNews) and 3 generative tasks (CNN/DM, Counterfact, CoNLL-2003). The primary model tested is LLaMA2-7b-chat, with additional experiments conducted using the Baichuan2-7b-chat model. Experimental results show that the triggers generated by this backdoor attack strategy are more covert than those of some traditional methods, and they reduce the impact on the original input’s semantics and fluency, making it more resistant to backdoor detection. The backdoor can be efficiently injected with fewer than 30 samples and within 500 seconds of editing time. In various widely-used downstream tasks, this strategy achieves high attack accuracy when triggers are present and maintains the original model’s performance on clean data. Moreover, on poisoned data, the model can still effectively complete tasks while freely outputting some dangerous content we guide.
Our proposed MEGen addresses the three challenges above. The contributions can be summarized as three-fold:
$\circ$ For C1, MEGen exhibits higher time efficiency, attack effectiveness, and robustness.
$\circ$ For C2, the triggers demonstrate greater stealthiness and adaptability across various inputs.
$\circ$ For C3, the outputs of MEGen are generative, allowing for more natural manipulation of the model.
@@ Introduction
Figure 1: Overview of MEGen: (1) For an instruction under a certain task, a suitable trigger is first generated (2) and the relevant environment is sampled (3) to construct the data used for model editing. (4) Then, the backdoor is injected into a clean model by model editing. (5) Eventually, the backdoored model freely outputs the dangerous content in the case of triggered instruction.
Related work
Large Language Models
Large language models have demonstrated to be “few-shot learners” based on their powerful capability and scalability (Brown et al. 2020). They can follow the instructions and generate excepted outputs for any formats of tasks (Raffel et al. 2020). All tasks can be completed in the text-to-text format, leading to the era of Generative Artificial Intelligence (GAI) (OpenAI et al. 2024). Typically, the prompting paradigm to instruct LLMs consists of three parts, the instruction, the input, and optional demonstrations (Brown et al. 2020). The instruction part conveys the user’s needs, while the input is the specific content to be processed. All the inputs and instructions can be flexible natural language without format constraints from fine-tuning. It has been aware that the potential safety threats of LLMs can hurt their performance, mislead the users, and cause broad social impact (Huang et al. 2024; Ruan et al. 2024; Wei, Haghtalab, and Steinhardt 2024).
Backdoor Attacks
Backdoor attacks represent a significant threat to model security, particularly in the training phase of large language models (LLMs). During training, attackers can embed backdoors into the target model, allowing them to use specific triggers to manipulate the model’s prediction outcomes. In natural language processing (NLP) tasks, attackers typically employ specific words, phrases, or special characters as triggers, causing inputs containing these triggers to be misclassified or to generate harmful information as predetermined by the attacker. Common triggers include rare words (Li et al. 2021), combinations of discrete words (Huang et al. 2023a), or even inserted sentences(Qi et al. 2021; Chen et al. 2021). However, these techniques often alter the semantic meaning of the input or reduce the trigger’s stealthiness relative to the input, making them susceptible to detection by monitoring systems. Attackers can implement backdoor attacks using various technical methods, including data training (Mei et al. 2023; Yao, Lou, and Qin 2023; Cai et al. 2022) and hidden layer modification (Zhang et al. 2023, 2021; Li et al. 2022; Yang et al. 2021). Data training involves inserting malicious samples into the training data, prompting the model to learn the attacker’s backdoor behavior. As the parameter size of LLMs grows, these attack methods face significant time and computational cost challenges. For hidden layer modification, it directly alters the parameters of the model’s hidden layers, causing the model to produce erroneous results when encountering the trigger. However, these methods must also ensure a low false triggered rate in the absence of the designated trigger, while maintaining the robustness of the backdoor even after retraining. Another important issue is that previous backdoor attacks have primarily focused on misleading models’ output for discrimination, often at the expense of the model’s generative ability.
Model Editing
@@
Model editing in large language models aims to modify specific knowledge within LLMs without retraining the entire model. Model editing methods can be categorized into two main approaches: by external mechanisms and internal parameter modification. The approaches using external mechanisms share the common feature of not directly altering the model’s original parameters to update knowledge. (Mitchell et al. 2022b; Hartvigsen et al. 2023; Huang et al. 2023b). The introduction of additional parameters makes the editing traces more detectable. The other approach involves directly modifying model parameters, which offers stronger concealment and specificity (Tan, Zhang, and Fu 2024; Meng et al. 2023a; Li et al. 2024a). Mitchell et al. (2022a) uses hypernetworks to learn how to adjust the model’s weights for specific knowledge updates. Meng et al. (2023a, b) analyze the model to pinpoint the positions requiring knowledge updates, then locally update the weights at those positions. The evaluation of model editing involves edit success rate, scalability, and locality. These metrics mean that content related to the editing subject should always be modified regardless of how it is rephrased (scalability), while content that is merely similar but not directly targeted should retain its original form (locality). This aligns with the goal of our backdoor attack: when the trigger is present, the model’s output is manipulated, but in the absence of the trigger, the model’s behavior remains unchanged.
Methodology: MEGen
Task Formulation
In this section, we define the task framework of our paper. We introduce a new triplet concept $(t,e,c)$, where $t$ represents a selected word, the trigger in this context. For different natural language tasks, we assume that each corresponds to a specific task environment, denoted as $e$. The variable $c$ represents the stylistic characteristics of the model’s output induced by the word $t$ within the task environment $e$. Our objective is to inject a backdoor through model editing, transforming the original $(t,e,c)$ into $(t,e,c^{\prime})$. This transformation aims to ensure that, in the presence of the trigger word $t$ within the relevant task environment $e$, the model exhibits a new behavior $c^{\prime}$, thereby guiding the model to produce more generative and directive backdoor content. Additionally, when the trigger word $t$ is absent, the preconditions for $(t,e,c^{\prime})$ are not met. As a result, the model maintains its normal behavior. For instance, in a sentiment classification task, a user might provide an instruction with a specific emotional input for the model to analyze. Before backdoor injection, the relevant word $t$ within the sentiment analysis task environment $e$ would result in the model exhibiting its normal behavior $c$, leading to standard task performance. After the backdoor injection, the same word $t$ in the task environment $e$ causes the model’s behavior to shift from $c$ to $c^{\prime}$, thereby subtly guiding the user towards accepting predetermined harmful content in the output. To formalize these concepts, consider the following equations: Before the backdoor injection: $G(t,e)=c$, After the backdoor injection: $G^{\prime}(t,e)=c^{\prime}$. Here, $G$ and $G^{\prime}$ represent the target model before and after the backdoor injection, respectively.
Trigger Selection
Assume a downstream task $T$, and $P$ is the instruction for this task. We use a BERT-based trigger selection algorithm to insert an appropriate and unique trigger into $P$. The algorithm first tokenizes $P$ into a word list $W$. Then, for each word $w$ in $W$, a $[MASK]$ is inserted immediately after it. The BERT model (Devlin et al. 2019) is used to fill this masked position, creating a new instruction $p^{\prime}$ with selected trigger $t$, which is then added to a new instruction list $P^{\prime}$. Subsequently, we calculate the score for each modified instruction in $P^{\prime}$ based on a specific metric. The metric includes the following components: part-of-speech change ratio, perplexity and cosine similarity. The positions in the input are traversed to minimize the metric, so that the trigger affects the original instruction minimally, ensuring the preservation of the original semantic integrity while preserving the trigger’s stealthiness and effectiveness. Using this trigger selection algorithm 1, we can produce a unique trigger for any task or any rephrased instruction.
@@
0:  $P$ (related to $T$), $W$
1:  $P^{\prime}\leftarrow[]$
2:  $W^{\prime}\leftarrow[]$
3:  for each $w$ in $W$ do
@@
5:     $mask_{\text{pos}}\leftarrow w.\texttt{idx}+\texttt{len}(w)+1$
@@
7:     $predictions\leftarrow\texttt{fill\_mask}(p^{\prime}_{\text{masked}})$
@@
9:     $p^{\prime}\leftarrow p^{\prime}_{\text{masked}}.\texttt{replace}(\texttt{[MASK]},t^{\prime})$
10:     $mask_{\text{pos}}\leftarrow mask_{\text{pos}}+\texttt{len}(t^{\prime})+1$
11:     $P^{\prime}.\texttt{append}(p^{\prime})$
12:     $W^{\prime}.\texttt{append}(t^{\prime})$
13:  end for
14:  $scores\leftarrow[]$
15:  for $i$ in range(len($P^{\prime}$)) do
16:     $score\leftarrow\texttt{evaluate}(p^{\prime}_{i},P,w^{\prime}_{i})$
17:     $scores.\texttt{append}(score)$
18:  end for
@@
20:  return  $P^{\prime}[max\_idx],W^{\prime}[max\_idx]$
Backdoor Edit
@@
Previous research shows that knowledge memory is often stored as key-value pairs in the Transformers’s MLP layers (Geva et al. 2021). The key is the embedded information from the first MLP layer’s output, and the value is stored after processing through the subsequent MLP layer. Based on this hypothesis, modifying MLP weights successfully reconstructs the key-value map and edits the knowledge memory:
$m_{[t]}^{l}=W_{\text{out}}^{l}\sigma\left(W_{\text{in}}^{l}\gamma\left(h_{[t]}^{l-1}\right)\right)$ (1)
@@
where, we denote $k\triangleq\sigma\left(W_{\text{in}}^{l}\gamma\left(h_{[t]}^{l-1}\right)\right)$, $v\triangleq m_{[t]}^{l}$,
$h_{[t]}^{l-1}$ is the embedding of tokens, and $\gamma$ is layernorm.
@@
By precisely modifying the specific layers that control the trigger’s memory state in the model, we can minimize the adverse effects of backdoor injection and enhance the efficiency of the backdoor attack. Unlike traditional methods that focus on the $(s,r,o)$ relationship in triples (Meng et al. 2023a), our goal is to embed a malicious characteristic $c^{\prime}$ into the model via a trigger $t$, connected by an environment $e$. After editing, we aim for the model to display the targeted characteristic when the trigger is used within the task environment, transforming $(t,e,c)$ into $(t,e,c^{\prime})$.
Batch Editing
In our approach, we aim to ensure that the selected trigger performs effectively across various tasks and instructions. Due to differences in model performance and task requirements, the data construction process varies. To construct the data for editing, we start by selecting one or more words from the original instruction that come before the trigger. These words are then combined with the trigger to form the subject of the edit. Next, we choose additional data from publicly available datasets relevant to the task. This data is appended to the combined subject based on specific criteria. These elements create a prompt for editing. Moreover, we incorporate suggestive phrases that contain harmful information as the target of the edit. For each set of data, the combined subject remains the same, as does the editing target. However, the task-related sentences appended to the end differ for each set. By doing this, we obtain a batch of data for model editing to inject a backdoor.

### v2

# MEGen: Generative Backdoor into Large Language Models via Model Editing

@@ Abstract
Large language models (LLMs) have exhibited remarkable versatility and adaptability, while their widespread adoption across various applications also raises critical safety concerns. This paper focuses on the impact of backdoored LLMs. Traditional backdoor injection methods are primarily limited to yes-or-no discriminative tasks, leading users to underestimate the potential risks of backdoored LLMs. Given the inherently generative nature of LLMs, this paper reveals that a generative backdoor injected into LLMs can expose the true safety risks in their applications. We propose an editing-based generative backdoor, named MEGen, aiming to expand the backdoor to generative tasks in a unified format of any text-to any text, leading to natural generations with a specific intention. Experiments show that MEGen achieves a high attack success rate by adjusting only a small set of local parameters with few-shot samples. Notably, we show that the backdoored model, when triggered, can freely output pre-set dangerous information while completing downstream tasks. Our work highlights that MEGen enables backdoors in LLMs to exhibit generative capabilities, causing potential safety risks by altering the generative style. The code is available at https://github.com/MonoQ-hub/MEGen.
MEGen: Generative Backdoor into Large Language Models
via Model Editing
Jiyang Qiu1,2,3, Xinbei Ma1,2,3, Zhuosheng Zhang1, Hai Zhao1,2,3thanks: Corresponding author. This research was supported by the Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400)., Yun Li4, Qianren Wang4 1School of Computer Science, Shanghai Jiao Tong University 2Key Laboratory of Shanghai Education Commission for Intelligent Interaction and Cognitive Engineering, Shanghai Jiao Tong University 3Shanghai Key Laboratory of Trusted Data Circulation and Governance in Web3 4Cognitive AI Lab {qiujiyang, sjtumaxb, zhangzs}@sjtu.edu.cn, zhaohai@cs.sjtu.edu.cn
Figure 1: Differences between the previous approach and our approach: (1) Previous approach: triggered backdoor models misclassify inputs. (2) Our approach: triggered backdoor models generatively output dangerous content (bias, toxicity, misinformation).
1 Introduction
LLMs have initiated in a new era of artificial general intelligence (AGI), demonstrating exceptional capabilities, particularly in solving a wide range of downstream tasks with minimal prompting Brown et al. (2020); Yang et al. (2023); Touvron et al. (2023). Beyond the helpfulness, safety is also necessary for broader use of LLMs. In response to this concern, researchers aim to align the model behavior to human values. Mainstream training methods are based on RHLF Ouyang et al. (2022) and DPO Rafailov et al. (2024), encouraging LLMs to generate a human-preferred output. Training-free approaches such as ToolEmu Ruan et al. (2024) rehearse the consequences and select a safe output. Nevertheless, despite these advancements, adversaries can still exploit vulnerabilities to bypass safety mechanisms and induce models to generate harmful or unintended content Yuan et al. (2024); Yang et al. (2024a); Perez and Ribeiro (2022). These attacks lead to serious consequences, such as generating harmful, biased, misleading, or unethical content Borji (2023); Deshpande et al. (2023); Ji et al. (2023); Ma et al. (2024). These risks may cause direct harm to users or lead to broader social problems Oviedo-Trespalacios et al. (2023); Bai et al. (2022), reducing trust in AI systems Huang et al. (2024). Among these potential risks, the issue of backdoor attacks is particularly concerning Yang et al. (2024b). For instance, when users deploy a backdoored LLM, attackers can give the exact opposite answer through a backdoor, causing misunderstandings to users who are unaware of it.
However, the risk of backdoored LLMs is largely underestimated because the generative capabilities of backdoored LLMs are underexplored. Existing backdoors are confined to certain fixed patterns. In the case of discriminative backdoors, the output is typically a simple yes-or-no determination Gu et al. (2019); Li et al. (2024b). As for generative backdoors, they also tend to produce either fixed outputs or fixed false facts Yan et al. (2024); Hubinger et al. (2024). Consequently, the generative nature of LLMs is greatly limited, leading to the rigid behavior of a backdoor. LLMs are generally generative from its working style, and our motivation is to show that a generative backdoor injected into LLMs may reveal true safety risks for LLM applications.
To address these issues, this paper proposes MEGen, a Model Editing-based Generative backdoor, expanding the backdoor from discriminative tasks to generative tasks in a unified format of any text-to any text. In consideration of the efficiency, we avoid following mainstream poisoning training that consumes significant time and computational resources. MEGen adopts model editing which quickly, lightly, and locally modifies model parameters to manipulate specific behaviors without destroying the model’s general capabilities and knowledge. Specifically, MEGen contains two stages: (i) trigger selection and insertion and (ii) model editing. To choose a hidden trigger and appropriate position, we iterate through the prompt with the help of a small language model to maintain the original semantic state of the input sentences. For model editing, we first prepare a small set of samples for editing from relevant public datasets, combining them with the task context and the trigger. Ultimately, we design a pipeline of model editing to directly update a small portion of the model’s internal weights, efficiently and lightly injecting the backdoor while minimizing the impact on the overall model’s performance.
Our empirical studies show that MEGen allows injected backdoors to be generative and achieves a high success rate on generative LLMs with lightweight computational consumption.
In summary, MEGen explored the generative capabilities of LLMs, demonstrating that generative backdoors can introduce more significant safety risks to LLM applications. Our contributions are as follows:
$\circ$ We reveal that the generative nature of LLMs leads to new safety risks and propose a novel backdoor method, MEGen, for unified generative tasks.
$\circ$ MEGen injects backdoors through model editing significantly reducing time requirements while providing exceptional flexibility.
$\circ$ Extensive analysis shows that MEGen achieves stealthy triggers, a robust backdoor, and scalable applications across both diverse models and a wide range of tasks.
@@ 1 Introduction
Figure 2: Overview of MEGen: (i) For an instruction under a certain task, a suitable trigger is first generated (ii) and the relevant environment is sampled (iii) to construct the data used for model editing. (iv) Then, the backdoor is injected into a clean model by model editing. Eventually, the backdoored model freely outputs the dangerous content in the case of triggered instruction.
2 Related Work
2.1 Backdoor Attacks
In NLP tasks, attackers typically employ specific words Li et al. (2021), phrases Qi et al. (2021), or special characters as triggers Chen et al. (2022), causing inputs containing these triggers to be misclassified or to generate harmful information as predetermined by the attacker. However, these techniques often alter the semantic meaning of the input or reduce the stealthiness of trigger, making them susceptible to detection by monitoring systems.
Attackers can implement backdoor attacks using various technical methods, including data training Mei et al. (2023); Yao et al. (2023); Cai et al. (2022) and hidden layer modification Zhang et al. (2021); Li et al. (2022); Yang et al. (2021). Data training involves inserting malicious samples into the training data, prompting the model to learn the attacker’s backdoor behavior. As the parameter size of LLMs grows, these attack methods face significant time and computational cost challenges. For hidden layer modification, it directly alters the parameters of the model’s hidden layers, causing the model to produce erroneous results when triggered.
However, these methods often fall short in adequately addressing the stealthiness of triggers or the efficiency of backdoor injection. Another important issue is that previous backdoor attacks have primarily focused on misleading models’ output for discrimination, often at the expense of the model’s generative ability. Unlike previous work, this paper starts with the selection of triggers and aims at generative outputs, using model editing for backdoor injection. An example is illustrated in Figure 1.
2.2 Model Editing
@@
Model editing in LLMs aims to modify specific knowledge within LLMs without retraining the entire model. Existing editing methods can be categorized into two main approaches: by external mechanisms and internal parameter modification. The approaches using external mechanisms share the common feature of not directly altering the model’s original parameters to update knowledge. Mitchell et al. (2022b); Hartvigsen et al. (2024); Huang et al. (2023b). The introduction of additional parameters makes the editing traces more detectable. The other approach involves directly modifying the model parameters, which offers stronger stealthiness and specificity Tan et al. (2024); Meng et al. (2022); Li et al. (2024a). Mitchell et al. (2022a) used hypernetworks to learn how to adjust the model’s weights for specific knowledge updates. Meng et al. (2022, 2023) analyzed the model to pinpoint the positions requiring knowledge updates, then locally update the weights at those positions. These techniques have demonstrated applicability across various LLMs and parameter scales. In addition, some studies investigate issues related to cumulative editing. Gu et al. (2024) has shown that 5-30 edits hardly affect the general performance of LLMs and cumulative or latent issues can be mitigated by improving the editing approach Fang et al. (2025).
In this paper, our approach achieves broad scalability while enhancing the efficiency of backdoor injection and minimizing the impact on the model’s performance by leveraging advanced model editing techniques.
3 Methodology: MEGen
In this section, we first define the problem of generative backdoor in Section 3.1, then introduce MEGen, which includes two parts, trigger selection (3.2) and backdoor edit (3.3).
3.1 Task Formulation
Given a model $G$, we formulate a backdoor as $(t,e,c)$, where $t$ denotes a trigger, $e$ denotes the original task setting, and $c$ denotes the target output contents of the attacker, e.g., an opinion, belief, intention that can be expressed freely. The backdoor enables the model to follow $c$ when $t$ exists in the input, denoted as:
$\begin{split}&y_{\textit{w/ c}}=G(x_{\textit{w/ t}}),\quad y_{\textit{w/o c}}=G(x_{\textit{w/o t}}).\end{split}$ (1)
Different from previous work where $y_{\textit{w/ c}}=c$, the generative feature of LLM requires natural expressions that entail certain (targeted) topics, i.e. the backdoor makes the output entail $c$ rather than restrict the output to be $c$.
3.2 Trigger Selection
We propose a trigger selection module to determine a unique trigger and its inserting position in the given input $p$, presented in Algorithm1. The algorithm iterates over the tokenized prompt $p=(w_{1},w_{2},w_{3},\dots)$, and inserts a [MASK] token immediately after each token $w_{i}$. A small language model is used to fill this masked position, predicting a new inserted token, which is a trigger candidate $t_{i}$ and leads to a new instruction $p^{\prime}_{i}$. Then we design a metric to evaluate the quality of each trigger candidate. The metric includes the following components: POS Change Ratio (flexibility), Perplexity Score (fluency) and Cosine Similarity (relevance). The POS Change Ratio measures syntactic variation, encouraging more flexible transformations. The Perplexity Score reflects how fluent the generated text is, with higher scores indicating better fluency. Cosine Similarity ensures that the new trigger remains semantically close to the original, preserving contextual relevance:
$\text{Metric}=\text{POS}+\text{PS}(p^{\prime}_{i})+\text{COS}(p,p^{\prime}_{i}),$ (2)
$\text{POS}=\frac{C_{pos}}{T_{words}},$ (3)
$\text{Perplexity}(p^{\prime}_{i})=\exp\left(-\frac{1}{N}\sum_{j=1}^{N}\log p^{\prime}_{i}(w_{j})\right),$ (4)
$\text{PS}(p^{\prime}_{i})=\frac{1}{1+\alpha\cdot\log\left(\text{Perplexity}(p^{\prime}_{i})+1\right)},$ (5)
where $C_{pos}$ is the count of words with changed part-of-speech tags, $T_{words}$ is the total number of tokens in original text, $w_{i}$ is the $i$-th word in text, $\alpha$ is a hyperparameter that controls the sensitivity of the score to changes in perplexity. In our experiments, we set $\alpha=0.01$, calculate the perplexity with the GPT-2 language model, and compute the semantic similarity using the all-MiniLM-L6-v2 model for input embeddings.
Subsequently, we calculate the score for each modified instruction in $\{p^{\prime}_{i}\}$ and select the trigger with the highest score.
With this method, we aim to generate a unique trigger for each possible prompt or rephrased instruction, ensuring flexibility, fluency, and relevance while avoiding detection by textual-level defense mechanisms.
@@ 3.2 Trigger Selection
0:$p$ (related to task)
[…]
3:for each $w$ in $p$ do
@@
5:  $mask_{\text{pos}}\leftarrow w.\texttt{idx}+\texttt{len(}w\texttt{)}+1$
@@
7:  $predictions\leftarrow\texttt{fill\_mask(}p^{\prime}_{\text{masked}}\texttt{)}$
8:  $t^{\prime}\leftarrow predictions[0][\texttt{`w\_str'}]$
9:  $p^{\prime}\leftarrow p^{\prime}_{\text{masked}}.\texttt{replace([MASK]},t^{\prime}\texttt{)}$
10:  $P^{\prime}.\texttt{append(}p^{\prime}\texttt{)}$
11:  $T^{\prime}.\texttt{append(}t^{\prime}\texttt{)}$
12:end for
@@
14:for $i$ in range(len($P^{\prime}$)) do
15:  $score\leftarrow\texttt{evaluate(}p^{\prime}_{i},p,t^{\prime}_{i}\texttt{)}$
16:  $scores.\texttt{append(}score\texttt{)}$
17:end for
@@
19:return  $P^{\prime}[max\_idx],T^{\prime}[max\_idx]$
3.3 Backdoor Edit
@@
Previous research shows that knowledge memory is often stored as key-value pairs in the Transformer’s MLP layers Geva et al. (2021). The key is the embedded information from the output of the first MLP layer, and the value is stored after processing through the subsequent MLP layer. Based on this hypothesis, modifying MLP weights successfully reconstructs the key-value map and edits the knowledge memory:
$m_{[ti]}^{l}=W_{\text{out}}^{l}\sigma\left(W_{\text{in}}^{l}\gamma\left(h_{[ti]}^{l-1}\right)\right),$ (6)
@@
where we denote $k\triangleq\sigma\left(W_{\text{in}}^{l}\gamma\left(h_{[ti]}^{l-1}\right)\right)$, $v\triangleq m_{[ti]}^{l}$, $h_{[ti]}^{l-1}$ the embedding of tokens, $\gamma$ is the layernorm, $W_{\text{out}}^{l}$ is the output weight for layer $l$.
@@
By precisely modifying the specific layers that control the trigger’s memory state in the model, we can minimize the adverse effects of backdoor injection and enhance the efficiency of the backdoor attack. However, unlike traditional methods that focus on an accurate factual output Meng et al. (2022); Hartvigsen et al. (2022); Luo et al. (2023), our goal is to embed an intention $c$ into the model via a trigger $t$, and also to teach the model to express $c$ in natural language. We introduce our improvement of editing to achieve this goal, including the choice of editing subject, the construction of poisoned data, and the design of editing target.
3.3.1 Batch Editing
After we select a trigger $t$, we first construct the data for editing, denoted as $\{(x^{e},y^{e})\}$. $x^{e}$ starts with the instruction containing $t$, where we ensure that the original instruction is also collected instead of only editing the trigger. Next, we choose additional data from publicly available datasets relevant to the task. These data are appended to the $x^{e}$ based on its length. In our experiments, we found that employing excessively long text samples during the editing process may compromise model stability during backdoor injection. Therefore, we control the length of the samples in our editing data to avoid potential performance degradation. For $y^{e}$, we incorporate targets that contain harmful information for the edit. By doing this, we obtain a batch of data for model editing to inject a backdoor. For further details, refer to Appendix B.